OpenStudy (anonymous):

If y varies inversely as x, and y = 5 as x = -2, find y for the x-value of 3? I said -10 or 10/3

OpenStudy (souvik):

y=k/x put y=5 and x=-2 find k

OpenStudy (anonymous):

-10 , I was right? or is it 5- -2?

OpenStudy (souvik):

right...now you get y=-10/x put x=3 and find y..

OpenStudy (anonymous):

-10/3 ! Thanks man, Helped alot!

OpenStudy (souvik):

my pleasure,,,:)